# ExploraCoder: Advancing code generation for multiple unseen APIs via planning and chained exploration

### v1

@@ 5.5 Ablation study
Table 7: Ablation study for GPT3.5 + ExploraCoder* on Torchdata-Manual.
@@
ExploraCoder* 8.76% 15.48% 14.04% 25.31% 16.75% 29.91% 19.81% 33.81%
w/o self-debug 4.48% 9.71% 8.71% 18.60% 11.61% 22.45% 13.79% 25.80%
w/o domain knowledge 3.90% 15.22% 6.86% 26.83% 7.90% 30.79% 9.82% 35.62%
w/o CoAE 0.85% 1.81% 3.38% 7.19% 5.33% 11.59% 7.30% 17.52%
w/o experience selection 5.81% 14.33% 9.53% 25.11% 9.98% 26.87% 10.00% 27.90%
We further conducted an ablation study on our best-performing framework, ExploraCoder*, in Table 7. We experiment on the Torchdata-Manual benchmark using GPT-3.5-turbo-0125. As discussed earlier, self-debug can effectively improve the performance of ExploraCoder. In practical deployment, users may consider incorporating more debugging iteration in CoAE to further boost the performance of ExploraCoder.
We then ablate the library-level domain knowledge (few-shot planner $\mathcal{D}$ and library introduction $s$) provided to ExploraCoder, and let the model plan API invocation subtasks based on its commonsense knowledge. We found that its execution success rate is comparable to that of ExploraCoder*, but the pass rate decreases. The model often breaks down problem into overly coarse-grained tasks that are difficult to solve, leading to buggy codes early in the chain. During the process of debugging and exploring along the chain, the LLM gradually drifts away from the original problem. Although it eventually produces executable code based on biased experience, the resulting program fails to fully meet the requirements.
We also ablate the CoAE by providing all the retrieved API information throughout ExploraCoder’s process to the generator, and prompt it to generate solution in a single run. We find that the performance significantly drop to 0.85 in pass@1. This suggests that (1) the current SOTA models still lack adequate in-context reasoning ability to handle complex tasks involving multiple API invocations, and (2) the information provided in API documentation is often insufficient, leading to hallucinations in API usage generation. This highlights the need for execution feedback to clarify ambiguities in API usage.
We further ablate a critical step within CoAE by removing the experience selection process. Instead of using the current selection strategy, we random select the candidates, ignoring their execution feedback. We find that the model still perform reasonably well on pass@1. However, as $k$ increases, its performance quickly converge to its limitation. We suggest this is because the model did not effectively leverage candidate experience information. An unexecutable API invocation experience provides limited insights to the API usage, and may even introduce some harmful effects to the final solution generation.
6 Conclusion
We present ExploraCoder, a novel code generation framework for LLMs to generate codes with multiple unseen API invocations through planning API invocation tasks and experiments with each task in a chain of API Exploration. We construct two challenging benchmarks Torchdata-Github and Torchdata-Manual to evaluate LLM’s performance when dealing with complex requirements that involve multiple unseen API invocation. Our experiments present the significant performance improvement of ExploraCoder when comparing with both related RAG frameworks and pretraining appraoches. We find that adding a self-debug component in the API exploration chain can further boost the performance of ExploraCoder. We provide comprehensive ablation studies to clearly present the benefits obtained by each of the components.
Appendix A Appendix
@@
A.1 Evaluating API knowledge retention in API-pretrained and API-untrained base models
@@
Listing 1: Demonstration of API knowledge in LLMs.
@@
--------- Output by CodeQwen1.5-7B-Chat ---------
@@
--------- Output by deepseek-coder-6.7b-instruct ---------
@@
A.2 Disccussion of fairness comparison between gpt-4-0613 and gpt-4-1106-preview.

### v2

@@ 5.5 Ablation study
Method $k=1$ $k=10$
Pass Success Pass Success
ExploraCoder* 11.5% 21.35% 20.87% 36.81%
w/o Self-Debug 7.00% 14.8% 13.84% 25.4%
w/o Lib-ICL 7.64% 15.73% 14.17% 29.77%
w/o CoAE 1.22% 2.21% 7.34% 13.38%
w/o Selection 4.12% 13.33% 9.46% 26.38%
@@
Table 4: Ablation study for ExploraCoder framework using GPT3.5 on Torchdata-Manual.
We further conducted an ablation study on our best-performing framework ExploraCoder* in Table 4. We experiment on the challenging Torchdata-Manual benchmark using GPT-3.5-turbo-0125.
As discussed earlier, self-debugging intermediate execution failure effectively improves ExploraCoder’s performance. This also suggests ExploraCoder may further benefit from dynamically generated testbed for intermediate code, which we will leave as exciting future work to explore.
We ablate the in-context learning of library-level knowledge (w/o Lib-ICL), removing the few-shot planner $\mathcal{D}$ and library introduction $s$, and let the model plan API invocation subtasks based soley on its commonsense knowledge. The performance decline observed could be attributed to the misalignment between planned subtasks and API granularity. Since overly coarse-grained subtasks introduce complexity, while incorrect subtasks that cannot be solved by any APIs increases the hallucination rates (Liu et al., 2024; Tian et al., 2024).
We ablate the CoAE (w/o CoAE) by providing all the retrieved API documentation throughout ExploraCoder’s process to the generator, and prompt it to end-to-end generate final solution. We find that the performance significantly drop to 1.22% in pass@1. This suggests (1) modern generators still lack adequate in-context reasoning ability to handle multiple unseen API invocations, and (2) API documentation could be insufficient, leading to hallucinated invocations. This highlights the need for intermediate execution or debug to gain more usage insights.
We further ablate a critical step within CoAE by removing the experience selection process (w/o selection). In this variant, candidate selection is randomized, disregarding executability signals. We find the success rate remains reasonably well, and the pass rate declines. A possible explanation is ExploraCoder degenerates into exploring low-quality API invocation chains with limited usage insights for fully accurate final solution.
6 Conclusion
We present ExploraCoder, a novel code generation framework for LLMs to generate multiple unseen API invocations through planning API-related subtasks and experimenting with each subtask in a novel chain-of-API-exploration. Experiments on our newly constructed benchmark and an existing benchmark demonstrates ExplroaCoder’s superior performance over competitive approaches.
7 Limitations
ExploraCoder’s effectiveness relies on the underlying LLM’s capabilities in handling long contexts and capturing API usage knowledge. Although our experiments show strong performance with both GPT-3.5 and GPT-4, small models with weak capability could exhibit less effectiveness on our complex multi-API-generation tasks. This dependency means that the framework’s performance is inherently bounded by the LLM’s capabilities. However, the rapid advancement in LLM development suggests this limitation may become less significant over time.
The framework assumes the availability of NL documentation, which may limit its effectiveness when dealing with overly incomplete, ambiguous, or erroneous API documentation. In our experiments, we simulated real-world scenarios by masking detailed parameter explanations and usage examples from the well-maintained torchdata documentation, approximating the minimal documentation typically available for newly introduced or privately maintained libraries. While this setting demonstrates ExploraCoder’s robustness with minimal API descriptions, future work could explore integrating additional knowledge sources, such as API client code or community discussions, to supplement insufficient documentation.
A promising improvement direction shared by ExploraCoder and related approaches is an early termination mechanism in the iterative generation workflow. When encountering particularly challenging problems where API exploration consistently fails, the system continues attempting solutions, potentially consuming unnecessary computational resources. The development of intelligent stopping criteria that can identify unsolvable problems or determine when further exploration would be unproductive represents an important direction for future research.
Appendix A Appendix
A.1 Comparing ExploraCoder with related approaches
We present a feature comparison in Table 5 and a detailed breakdown of the computational overhead (model call and token consumption) in Table 6.
A.1.1 Feature Comparison
ExploraCoder introduces two key innovations in LLM-based code generation: exploratory planning and step-wise code construction. (1) Firstly, traditional approaches generally follow either "preactive" planning (based on prior knowledge) or "reactive" planning (based on environmental feedback). Preactive planning, such as CoT prompting, can suffer from hallucinations when handling complex or out-of-distribution tasks. Reactive planning, common in agent-based systems, often lacks systematic consideration and controllability (Xia et al., 2024). ExploraCoder bridges this gap by introducing exploratory planning, which enhances preactive plans with step-wise environmental feedback to mitigate hallucinations while maintaining systematic control. (2) Secondly, most existing work conduct end-to-end code generations/modification. We propose a step-wise code construction workflow to generate partial code for a simple subtask based on the planning instructions. Le et al. (2024) exhibits a similar idea of preactively planning a series of reusable functions for simple self-contained code generation, while it fails to leverage the step-wise execution signal from partial codes and it is also not applicable to more complex programming scenario like multi-unseen-API invocations.
ExploraCoder vs. Existing library-oriented approaches.
While existing library-oriented approaches (DocPrompting, CAPIR, EpiGen) primarily focus on API-docs retrieval quality, ExploraCoder addresses the fundamental limitations in LLMs’ multi-API reasoning capabilities and documentation ambiguity. Through a novel Chain of API Exploration, ExploraCoder iteratively collects execution information to resolve API usage ambiguities. For instance, in task_175, traditional preactive planning in EpiGen hallucinated about the parameter types in "LineReader". ExploraCoder resolves such issues by executing multiple candidate implementations for the "read the lines" subtask and filtering out incorrect API usage patterns, thereby acquiring additional API knowledge that cannot be derived from documentation alone.
Method API Retrival Planning Step-wise Code Construction Debugging Manual Requirement
DocPrompting -
CAPIR Preactive -
EpiGen Preactive -
ExploraCoder Exploratory -
Self-Repair Reactive -
ReAct Reactive Agentic traj. fewshot
KnowAgent Reactive Agentic traj. fewshot
CAPIR + Self-Repair Reactive -
ExploraCoder* Exploratory -
Table 5: Features in each retrieval-based baselines. $\dagger$: CAPIR focuses exclusively on planning for the retrieval phase, without addressing the generation phase.
Model Pre-Processing Calls Code Generation Calls Overall Model Calls Tokens Pass@10
DocPrompting 0 1 1 10k 8.15%
CAPIR n+2 1 3+n 18k 11.23%
EpiGen n+2 1 3+n 18k 13.40%
ExploraCoder n+2 n+1 3+2n 56k 17.76%
CAPIR + Self-Repair n+2 1.5n+2 4+2.5n 70k 15.19%
ExploraCoder* n+2 2.6n+1 3+3.6n 95k 24.26%
ReAct - - 2N+2 112k 7.43%
KnowAgent - - 3N+3 143k 13.51%
Table 6: Computational costs and performance in each retrieval-based baselines.
ExploraCoder vs. iterative-debugging.
ExploraCoder’s step-wise code construction offers significant advantages over existing iterative-debugging approaches. While methods like CAPIR+Self-Repair employ bug-driven reactive planning on complete code solutions, ExploraCoder* debugs simpler subtasks at earlier stages, preventing the accumulation of complex errors. We observe in case study (Appendix A.9) that CAPIR+Self-Repair repeatedly attempts to fix a buggy code that deviates substantially from the correct solution, continuing until it exhausts its iteration budget.
ExploraCoder vs. Agent-style frameworks.
ExploraCoder differs from agent-style frameworks in three crucial aspects:
1) Structured Workflow: Unlike agent frameworks with undeterministic actions, ExploraCoder implements a pinned step-wise code construction workflow (exploratory programming) that experimentally performs well with multi-API invocations. Most code agents, eg. Swe-Agent (Yang et al., 2024), MetaGPT (Hong et al., 2024), conduct end-to-end code construction/modification, just like what we discussed with Self-Repair. In our agentic baseline implementations (Appendix A.1.3), we borrow the idea of ExploraCoder and prompt React and KnowAgent to generate partial code at each step. Our empirical results show that enforcement of step-wise code generation in agentic workflow is often unstable and uncontrollable, which aligns with the suggestions with Agentless (Xia et al., 2024).
2) Systematic Planning: ExploraCoder’s exploratory planning maintains a comprehensive view of task dependencies, preventing common pitfalls seen in reactive planning. For example, in task_124 (CSV loading from compressed files), reactive agents often overlook crucial steps like decompression, leading to inefficient API retrieval cycles. ExploraCoder’s exploratory planning first systematically breaks down such tasks into logical, dependent steps. Then it also leverages the environmental feedback to enhance the next-step plannings with additional API knowledge.
3) Efficiency and Accessibility: Agent-based approaches require high-quality few-shot examples for reasoning trajectories, which are often impractical when working with new libraries. We empirically find ReAct and KnowAgent performance deteriorize when we remove the examplary trajectory or provide an OOD trajectory on other libraries (See implementation details in Appendix A.1.3). Additionally, they tend to be token-inefficient due to potential deterioration into recursive or meaningless actions when facing noisy observations.
A.1.2 Computational Comparison
From Table 6 we can observe that ExploraCoder is cost-efficient. When Compared to CAPIR+Self-Repair and two agentic methods, ExploraCoder achieves higher performance with fewer model calls and lower token consumption. Additionally, ExploraCoder* achieves a 59.7% performance improvement over CAPIR+Self-Repair with a 35.71% increase in token consumption and a 44% increase in model calls, demonstrating that the performance gains significantly outweigh the proportional increase in resource consumption.
Calculation details: For model calls, we provide clear analytical expressions based on n, the number of decomposed subtasks. For token consumption, we randomly sampled a task subset with n=8 (the mean value of n in our datasets), and generated 20 candidate final solutions, then calculated the average token consumption per task. Note that the agent-based methods’ action is uncontrollable, and their model calls cannot be mapped to a preactively determined n, therefore we use a different N to represent its iteration steps. We empirically observe N>n in most tasks. For the two non-agent approaches involving self-debug mechanisms, Self-Repair and ExploraCoder*, debugging rounds is not deterministic. Therefore we use the formulated expectation based on the empirically observed debug rate in our experiments. Having the probability of the two methods conducting debug as $p_{1}$ (ExploraCoder*) and $p_{2}$ (Self-repair), their expectation of model call can be formulated as $(1+5p_{1})n+1$ and $22p_{2}n+$ 2. Notably, $p_{2}=0.75$ is significantly higher than $p_{1}=0.32$ across two benchmark. This debug rate difference arises because ExploraCoder* focuses on debugging simple intermediate subtasks (which are generally less error-prone), while Self-Repair always attempts to debug a complete solutions (which often fail to repair successfully, triggering additional debugging iterations up to the budget limit).
A.1.3 Implementation details of KnowAgent and ReAct
We use the official code and prompt released to implement ReAct (Yao et al., 2022) and KnowAgent (Zhu et al., 2024):
Action Space:
In our experiment, we abstract the capability of ExploraCoder* and design the following actions/tools for ReAct and KnowAgent:
  1. 1.
    Retrieve[target_functionality]: Query Torchdata API documentation for a specific functionality, returning top-k relevant APIs.
  2. 2.
    Write_and_Execute[code]: Generate/Debug then execute an in-progress code snippet to test partial functionality. The execution information will be returned.
  3. 3.
    Finish[code]: Write the complete code solution that solves the coding task based on reasoning trajectroy.
Trajectory example:
We manually crafted one long trajectory example of the agent-style Torchdata task-solving process across 8 reasoning steps, showing the interleaving trace of API retrieval and code generation, code debug, and more API retrieval, etc. While this enables the agent to understand the expected reasoning flow, we note this manual involvement is expensive in real-world deployment, especially for newly-introduced libraries. We empirically observe their performance deteriorize when we remove the examplary trajectory or provide an OOD trajectory on other libraries.
Reasoning steps & model call budget:
To ensure fair comparison given our tasks’ complexity, we extended the reasoning step budget of React and KnowAgent to 16 on Torchdata-Github and 21 on Torchdata-Manual (vs. original 10), enabling them to initiate 32/42 and 48/63 model calls to perform analysis, planning, and write/debug code snippets. Table 6 shows that the token consumption are significantly higher than non-agent baselines.
(a) success-fail division
for final solution in
ExploraCoder
(b) pass-fail division
for final solution in
ExploraCoder
(c) success-fail division
for final solution in
ExploraCoder*
(d) pass-fail division
for final solution in
ExploraCoder*
A.2 Comparing ExploraCoder with repeated-sampling
Inferior baselines sometimes have simpler framework designs thus costing fewer tokens. Their performance could naturally be improved through scaling up the inference-time computation in the sampling phase. We investigate whether repeated-sampling token-efficient methods like CAPIR can achieve competitive performance comparable to token-intensive approaches like ExploraCoder* with equivalent token budgets. Experiments on 50 randomly sampled problems from Torchdata-Manual in Table 7 shows that:
Method Basic Metrics Extended Metrics
Pass@10 Pass@20 Pass@50 Pass@90 Pass@100 Pass@120 Pass@130
CAPIR 8.48% 9.70% 11.84% 13.33% 13.52% 13.71% 13.72%
ExploraCoder 16.07% 17.55% not evaluated
ExploraCoder* 24.99% 27.22% not evaluated
Table 7: Pass@k performance comparison.
1) Efficient method like CAPIR doesn’t neccessarily yield competitive results by repeated-sampling.
Under same level of token consumption  [^3] Token calculation details: As shown in Table 7, the fixed token difference between CAPIR and ExploraCoder is around 38k($\approx$ \$0.038), let a marginal token consumption per sample be 450 (empirically on Torchdata-Manual), we can equalize the token consumption between CAPIR and ExploraCoder when CAPIR have 84 more samples than ExploraCoder, which means token(CAPIR, k=100) $\approx$ token(ExploraCoder, k=20) , CAPIR k=100 (pass=13.52%) underperform than ExploraCoder k=20 (pass=17.55%). And it even underperfrorms than ExploraCoder with k=10 (pass=16.07%).
2) The improvement in pass rate plateaus as the sampling number k increases.
CAPIR’s pass rate barely improve when we scale the k from 100 to 130. This aligns with Brown et al. (2024)’s findings that inference-time sampling typically follows logarithmic scaling laws. This indicate inferior methods like CAPIR should take even much larger computational costs to possibly achieve adequate performance with ExploraCoder.
3) Large repeated-sampling number present fundamental fairness and practicality concerns.
While evaluating methods under equivalent computational budgets provides valuable insight, we’d like to suggest that raising large sampling numbers (eg. $\geq$100) creates practical challenges for candidate verification, which limits its real-world applicability. Brown et al. (2024) indicates that despite the effectiveness of repeated sampling for correct answer, it is often hard for users (verifiers) to verify a candidate solution from the large volumes of samples. For example, test suites are often inaccessible during development of new functionalities (Siddiq et al., 2023). Furthermore, managing a large number of candidate solutions does not align with typical development workflows.
Our experiments demonstrate that ExploraCoder standsout in two critical aspects: (1) its ability to generate correct solutions with a manageable sampling budget, and (2) its superior token efficiency compared to computation-scaling variants of CAPIR (including both w/ repeated-sampling and w/ Self-Repair workflow). These results highlight ExploraCoder’s practical advantages in real-world applications, where both solution quality and computational efficiency are essential considerations.
A.3 Quantitative analysis
A.3.1 The effectiveness of CoAE in ExploraCoder
ExploraCoder leverages API invocation experience from CoAE to enhance the quality of final solution generation. Intuitively, the quality of exploration subtasks within CoAE is closely related to the quality of the final solutions.
To explore the pass/success rate relationship between CoAE subtasks and final solutions in ExploraCoder, we conducted a quantitative analysis, examining how the number of CoAE subtasks and their success rates affect the pass rate and overall success rate of the final solutions. We illustrate their correlation using results from the best-performing base model, GPT-4-0613, on our Torchdata-Manual benchmark.
From Figure 3(a) to 3(d), we observe that both the pass rate and success rate of the final solutions positively correlate with the CoAE subtask success rate. Subtasks with higher success rates, particularly those with a success rate of 1, are more likely to generate successful or passing final solutions. Interestingly, the number of subtasks doesn’t appear to have a significant direct impact. However, as shown in Figures 3(a) and 3(b), without intermediate self-debug, problems with a higher subtask number (ranging from 10 to 13) tend to have lower subtask success rate as the subtask number increase. This may be due to the increased complexity of inter-task API interactions. When the self-debug mechanism is introduced in ExploraCoder*, we observe in Figures 3(c) and 3(d) a notable improvement in the overall subtask success rate, even for cases with higher subtask numbers. This leads to more successful and passing final solutions. The improvement can be attributed to ExploraCoder’s ability to correct typos and simple API interaction errors in each subtask, thereby gaining richer API usage experience and exploiting it to the final solution generation.
A.3.2 The effectiveness of task planning in ExploraCoder
#API #decomposed subtask API per subtask
Torchdata-Github 4.26 4.06 1.04
Torchdata-Manual 9.94 8.22 1.21
Table 8: Summary of decomposed subtask statistics.
A.4 Construction details of Torchdata-Manual
Benchmarks
Num.
samples
Num.
APIs
Num.
Invoc.
Volume of
doc pool
Torchdata-Github 50 3-8 3-8 228
Torchdata-Manual 100 8-14 8-21 228
Table 9: Statistical Summary of two Torchdata-based benchmarks. Num. APIs reports the range of distinct APIs involved in each sample. Num. Invoc. reports the range of API invocations in the samples’ canonical solution. Volume of the doc pool refers to the number of API documents provided by the library, which also represents the size of the search space during API retrieval.
The Torchdata-Manual benchmark is designed to provide complex programming problems that require the use of multiple Torchdata APIs. It follows the style of prior unseen library benchamarks (Zan et al., 2023; Ma et al., 2024) , consisting of a natural language task description, code context, canonical solutions, and test cases. The construction process is outlined as follows:
Expert Review. Finally, we invited two Python programmers, each with four years of experience, to review the dataset and suggest adjustments. Specifically, we ask the experts to examine on 4 aspect of the crafted programming problems (1) The executability of the canonical solution, (2) The intuitiveness of the API usage, (3) The rigor of the test cases, (3) The meaningfulness of the task requirements. If any issues were identified in these aspects, the experts discussed them with the task creators and revised the tasks accordingly. This step ensured the overall quality and correctness of the benchmark. All participants were compensated adequately, with payment aligned to ethical standards and appropriate to their demographic and region. And we ensure that there is no ethical issue involved in our data construction process.
A.5 Generalizabiliy of ExploraCoder on More Unseen Libraries Settings
@@
Direct Generation 3.40% 3.89% 6.67% 8.34% 7.80% 10.56% 8.16% 12.15%
DocPrompting (2023) 8.00% 21.61% 14.01% 40.85% 18.30% 48.84% 19.70% 54.30%
EpiGen (2024) 7.77% 23.13% 14.41% 36.47% 17.86% 41.47% 20.30% 46.54%
ExploraCoder (Ours) 19.73% 59.47% 25.75% 67.36% 27.36% 68.76% 28.47% 69.39%
CAPIR + Self-Repair (2024) 13.90% 33.14% 20.78% 50.95% 23.64% 56.59% 24.49% 59.09%
ExploraCoder* (Ours) 21.33% 74.87% 26.73% 79.50% 28.87% 80.07% 30.39% 80.56%
Table 10: Comparing ExploraCoder with advanced retrieval-based approaches using GPT3.5 on MonkBeatEval.
While Torchdata has become an established practice for evaluating unseen library code generation (Zan et al., 2023; Zhang et al., 2023; Ma et al., 2024), it is important to assess ExploraCoder’s effectiveness across a broader range of libraries. However, as noted by Zan et al. (2023), it is particularly challenging to discover a suitable library like TorchData in open-source communities.
To enable rigorous evaluation of ExploraCoder’s generalizability, we introduce MonkBeatEval, a new multi-library benchmark constructed specifically for testing performance on unseen libraries.
A.5.1 Construction of MonkBeatEval
Library Creation. We developed two pseudo-libraries (Monkey and BeatNum) by applying bidirectional transformation mappings to Pandas and NumPy APIs, using the methodology [^4] Zan provide in their paper a set of transformations rules between Pandas/Numpy and Monkey/BeatNum from Zan et al. (2023), eg. "pandas.iterrows"$\leftrightarrow$"monkey.traversal”. This ensures the APIs are novel to LLMs.
Programming Problem Construction. We adapted 50 multi-API problems from PanNumEval (Zan et al., 2024), applying the same transformations to create semantically equivalent tasks using Monkey and BeatNum APIs. We carefully examine the transformation result and extend the transformation rules where some library information was not converted. Furthermore, we masked out the explicit API usage in task description like “… using np.linspace” to prevent unintentional hints about the original libraries.
Execution Framework. We implemented real-time bidirectional API transformation to enable execution-based evaluation using pseudo libraries, while maintaining the illusion of two new libraries for LLM’s interactions. Specifically,
  1. 1.
    For executing the LLM generated code snippets: Pseudo API calls→Original API calls
  2. 2.
    For LLM obtaining the intermediate output&exceptions: Original info $\rightarrow$ Pseudo info
A.5.2 Evaluation Results on MonkBeatEval
We replicated representative related work that exhibited most competitive performance in Section 5.4 on MonkBeatEval using GPT-3.5. The results are presented in Table 10.
The trends observed in the results are consistent with the discussions in Section 5, with ExploraCoder achieving SOTA performance. This consistency further demonstrates the robustness of our evaluation, showcasing ExploraCoder’s effectiveness and generalizability across different library contexts and integration scenarios.
@@ A.5.2 Evaluation Results on MonkBeatEval
A.6 Evaluating prior API knowledge retention in API-pretrained and API-untrained base models
@@
Listing 1: Demonstration of prior API knowledge in different LLMs.
@@
----------------------------------------------- Output by CodeQwen1.5-7B-Chat -----------------------------------------------
@@
------------------------------------------- Output by deepseek-coder-6.7b-instruct -------------------------------------------
@@
A.7 Disccussion of fairness comparison between gpt-4-0613 and gpt-4-1106-preview.